OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 4) and (−2, 13)? 3x + y = 7 3x − y = 7 x − 3y = 7 x + 3y = 7

OpenStudy (anonymous):

(y2-y1)/(x2-x1) = slope

OpenStudy (anonymous):

(13-4)/(-2-1)=-3

OpenStudy (anonymous):

Then use point slope form (y-y1)=m(x-x1)

OpenStudy (anonymous):

y-4=-3(x-1)

OpenStudy (anonymous):

y-4=-3x+3

OpenStudy (anonymous):

y=-3x+7